A diffraction grating has 5500 lines/cm . When a beam of monochromatic light goes through it, the second pair of bright spots occurs at ±26.0 cm from the central spot on a screen 43.0 cm past the grating.

A. What is the wavelength of this light?

B. How far from the central spot does the next pair of bright spots occur on the screen?
Answer #1

d = 1/5500/cm

d = 1818 nm

d sintheta = m*lambda

tan theta = y/D

theta is small tan theta = sin theta

d*y/D = m*lambda

lambda = d*y/D*m

y = 26 cm

D = 43 cm

m = 2

lambda = 549.68 nm

part b )

d*y/D = m*lambda

m =3

y = m*lambda*D/d

y = 39 cm